OpenStudy (anonymous):

if a regular can is 12 cm tall and sells for $1.25 and a large can is the same shape and 16cm tall, and sells for $2.00, how much more does the large can hold?

OpenStudy (campbell_st):

do a simple comparison let the radius = 1 unit Volume of a Cylinder is \[V = \pi r^2h\] so \[V = \pi 1^2 \times 12.... and V = \pi 1^2 \times 16\] so the volume is increasing by \[16\pi - 12\pi = 4\pi\] you can see if this is still the case when the radius is 2 units

OpenStudy (anonymous):

oops, I meant to say how many more TIMES does the larger can hold than the smaller.

OpenStudy (anonymous):

I am sorry to seem like an idiot, but I am not sure how I am to get the exact answer if I am using a radius that was not given in the problem?

OpenStudy (anonymous):

\[\frac{\text{Vlarge}}{\text{Vregular}}=\frac{2.}{1.25} \]\[\text{Vlarge}=1.6 \text{ * Vregular} \]